OpenStudy (anonymous):

simplify (5 sin 6α)/ (cos^ 2 3α − sin^ 2 3α)

myininaya (myininaya):

\[\frac{5 \sin(6 x)}{\cos(2 \cdot 3x)}=\frac{5 \sin(6x)}{\cos(6x)}=5 \tan(6x)\]